Q Search this course Homework of 70 n ±8 of the population mean (to 4 decimals)? Check
TABLE 1 CUMULATIVE PROBABILITIES FOR THE STANDARD NORMAL DISTRIBUTION (Continued) probability Entries in the table give the area under the curve to the left of the z valuc. For example, for z-1.25, the cumulative probability is 8944 0 5000 5040 5080 5120 5160 5199 5239 52795319 5359 1 5398 5438 5478 5517 55575596 5636 5675 5714 5753 2 5793 5832 5871 5910 59485987 6026 6064 6103 6141 3 6179 6217 6255 6293 6331 6368 6406 64436480 6517 4 6554 6591 6628 6664 67006736 6772 6808 68446879 5 6915 6 6985 7019 7054 70887123 7157 7190 7224 6 7257 7291 7324 7357 7389 7422 7454 7486 7517 .7549 7 7580 7611 7642 7673 7704 77347764 7794 78237852 8 .7881 7910 7939 967 7995 80238051 80788106 8133 9 8159 8186 821282388264 8289315 83408365 8389 10 8413 8438 8461 8485 8508 85318554 8577 8599 862 1.1 8643 8665 8686 8708 829 8749 87708790 88108830 1.2 8849 8869 .8888 8907 8925 89448962 89808979015 13 9032 9049 9066 9082 9099 9115 9131 9147 9162 9177 1.4 .9192 .9207 .9222 .9236 .9251 9265 .9279 9292 .9306 .9319 15 9332 9345 9357 9370 93829394 9406 9418 9429 944 1.6 9452 9463 9474 9484 9495 95059515 9525 9535 9545 1.7 9554 9564 9573 9582 95919599 96089616 9625 9633 18 9641 9649 9656 9664 967196789686 9693 9699 906 1.9 97139719 9726 9732 9738 974497509756 9761 9767 20 9772 9778 9783 9788 9793 97989803 9808 9812 9817 21 .9821 9826 .9830 .9834 .9838 .9842 .9846 .9850 .9854 .9857 22 9861 9864 98689871 987598789881 9884 9887 9890 23 9893 98969898 9901 9904906 990919913 9916 24 9918 9920 9922 9925992799299939932 9934 9936 25 993899409941 994399459946 994899499951 9952 2.6 9953 995599569957 995999609961 996299639964 2.7 9965 9966 9967 96899699970997997299739974 2.8 9974 9975 997699799779978 979 9979 9580998 29 9981 9982 9982998399849984 998S 9985 9986996 3.0 99879987 9987 99889988 99899989 9989 9990 9990
Answer #1

Answer:

a).

standard error = sd/sqrt(n) = 70/sqrt(125) =6.2610

we have to find P( 292 <mean x < 308)

z value for 292, z =(292-300)/6.2610 = -1.28

z value for 308, z =(308-300)/6.2610 = 1.28

P( 292 <mean x < 308) = P( -1.28 <z<1.28)

=P( z < 1.28)-P( z < -1.28)

= 0.8997 - 0.1003

=0.7994

b).

we have to find P( 289 <mean x < 311)

z value for 289, z =(289-300)/6.2610 = -1.76

z value for 311, z =(311-300)/6.2610 = 1.76

P( 289 <mean x < 311) = P( -1.76 <z<1.76)

=P( z < 1.76)-P( z < -1.76)

= 0.9608-0.0392

=0.9216